# Bayesian dictionary learning estimation of cell membrane permeability from surface pH data

### v1

@@ 3.1 Bayesian sparsity promoting linear least squares solvers
$\pi_{X\mid\Theta}(x\mid\theta)\sim\pi_{+}(x){\rm exp}\left(-\frac{1}{2}\sum_{j=1}^{p}\log\theta_{j}-\frac{1}{2}\big{\|}{\mathsf{D}}_{\theta}^{-1/2}x\big{\|}^{2}\right),$
@@
[…]
@@ 5 Computed example
[…]
@@
To identify the best explaining subdictionary for each of the test cases, we run the IAS algorithm with the sparsity parameter $\eta=0.001$, and the stopping criterion for the outer iteration $\|\theta^{k+1}-\theta^{k}\|<\text{tol}_{\theta}=0.001$. The maximum number of IAS iterations allowed is 1 000.
@@
[…]
@@
Experiment 1 Experiment 2 Experiment 3 True Est True Est True Est $\lambda$ $[\mu{\rm m}/{\rm s}]$ 30.78 30.84 27.36 27.43 27.36 27.29 $A^{0}$ 16 16.29 20.0 19.36 16.0 17.31 $\gamma$ $[\mu{\rm m}/{\rm s}]$ $10^{-4.2}$ $10^{-4.19}$ $10^{-4.0}$ $10^{-4.01}$ $10^{-4.15}$ $10^{-4.12}$
@@ 6 Conclusions
In this article, we propose a layered dictionary matching algorithm for the estimation of model parameters of a computational model describing the pH response due to gas transport through a cell membrane in a presence of a measurement device that itself has an effect on the microenvironment. The main characteristics of this model is that while the parameter space is low-dimensional, the forward model is complex and computationally intensive. Moreover, the non-linearity of the forward model makes the analysis of optimization-based parameter estimation algorithms difficult, while, e.g., sampling-based solutions of the inverse problem may be too time consuming. In [3], the authors considered the inverse problem in the framework of Bayesian particle filtering techniques which avoids the time-consuming forward solutions of the full model over the experimental time window, requiring only local solutions over short time steps. Compared to the particle filtering method with $4800$ particles and $1001$ measurements at $2{\rm Hz}$, the computing times of the proposed dictionary learning method is, given similar hardware, approximately one fifth. This comparison takes into account the preprocessing time comprising the creation of the dictionary, clustering, extraction of the features and dictionary compression error estimation. We point out that once the preprocessing phase has been completed, solution of the inverse problem is very fast, thus improving the performance by orders of magnitude.
@@
The technique proposed in this paper appears useful in the case in which the estimation of parameters must be carried out with limited computational resources, in a relatively short time, or when several estimates must be run in parallel starting with the same dictionary. The obtained estimates appear to be relatively accurate, in fact, the predicted outputs with estimated parameters are indistinguishable from the original data, as Figure 4 suggests. A limitation of the proposed algorithm is its reliance on dense sampling of the parameter space which was low-dimensional in the present example. In high-dimensional parameter models, the parameter space sampling requires further thoughts, e.g., by relying multilevel sampling or Monte Carlo sampling with some extra prior information that guides the sampling process.

### v2

@@ 3.1 Bayesian sparsity promoting linear least squares solvers
$\pi_{X\mid\Theta}(x\mid\theta)\propto\pi_{+}(x){\rm exp}\left(-\frac{1}{2}\sum_{j=1}^{p}\log\theta_{j}-\frac{1}{2}\big\|{\mathsf{D}}_{\theta}^{-1/2}x\big\|^{2}\right),$
@@
  • Given: matrix ${\mathsf{A}}\in{\mathbb{R}}^{m\times p}$, $b\in{\mathbb{R}}^{m}$, $\beta$, $\vartheta_{j},\;1\leq j\leq p$,
  • Initialize: $\theta^{(0)}_{j}=\vartheta_{j},\;1\leq j\leq p$, set the counter $k=0$.
@@
The IAS algorithm has been proved to be flexible and easy to modify to accommodate the needs of the problem in question, as it reduces the sparsity promoting inversion in two technically simple steps, a least squares problem with possible constraints for updating $x$, and a function evolution for updating $\theta$. The formulation of the first step as a standard least squares problem makes it possible to take advantage of the power of numerical linear algebra for the solution of linear least squares problems. Krylov subspace least squares solvers are the methods of choice for very large problems, including those in which the matrix cannot be stored in the memory, or when the linear operator is defined in a matrix-free form.
@@ 5 Computed example
To estimate the dictionary compression error, we fix the sample size of the compression errors to $K_{i}=7\,000$ for each subdictionary. For each subdictionary, we draw from the regular grid in the parameter space a random parameter vector $\widehat{\xi}$ that is associated to the $i$th dictionary. To generate a data vector that presumably represents the $i$th dictionary, we perturb the components of $\widehat{\xi}$,
@@
To simulate the limited measurement precision, the data was rounded off with the precision of 0.02 pH units. To identify the best explaining subdictionary for each of the test cases, we run the IAS algorithm with the sparsity parameter $\eta=0.001$, and the stopping criterion for the outer iteration $\|\theta^{k+1}-\theta^{k}\|<\text{tol}_{\theta}=0.001$. The maximum number of IAS iterations allowed is 1 000.
@@
the right hand side being equal to the expected square of whitened noise with T=1 001T=1\,001 independent components.
@@
Experiment 1 Experiment 2 Experiment 3 True Est True Est True Est $\lambda$ $\mu{\rm m}/{\rm s}]$ 30.78 30.84 27.36 27.43 27.36 27.29 $A^{0}$ 16 16.29 20.0 19.36 16.0 17.31 $\gamma$ $\mu{\rm m}/{\rm s}]$ $10^{-4.2}$ $10^{-4.19}$ $10^{-4.0}$ $10^{-4.01}$ $10^{-4.15}$ $10^{-4.12}$
@@
Finally, we compare the computingtimes of the proposed algorithm and the particle filter algorithm [3]. While both algorithms can be run without problems in a personal computer environment, for the timings of the test runs reported below the CINECA Leonardo HPC environment was used. In Table 6, shows the timing for each subtask of the dictionary matching algorithm is listed separately. The full calculation using the dictionary of size 64000 atoms from end-to-end takes less than 2 and a half hours, and with precomputed dictionary, the matching process requires about three minutes. For comparison, the parameter estimation using the particle filtering approach with 4800 particles on the same hardware requires from 4 hours 45 minutes to 5 hours using 112 cores, or from 7 hours 30 minutes to 8 hours using 56 cores. The comparison shows that even if the full dictionary is created from scratch, the computing time is reduced by a factor of two or more. Using a precomputed dictionary, the speedup from hours to few minutes is dramatic.
Task # cores time Dictionary creation 40 1 hour 10 minutes Distance matrix 112 33 minutes $k$-medoid clustering 112 << 10 minutes Rank estimation 112 seconds NMF factorization 112 4 minutes Error density estimation 112 23 minutes Dictionary matching 8 3 minutes
Table 6: Timing of the subtasks of the dictionary matching algorithm with a dictionary comprising 64000 atoms. The computations were performed on the CINECA Leonardo HPC having 112 cores/node: Each node has two Intel Sapphire Rapids Intel Xeon Platinum 8480+ processors of 56 cores.
@@ 6 Conclusions
In this article, we propose a layered dictionary matching algorithm for the estimation of model parameters of a computational model describing the pH response due to gas transport through a cell membrane in a presence of a measurement device that itself has an effect on the microenvironment. The main characteristics of this model is that while the parameter space is low-dimensional, the forward model is complex and computationally intensive. Moreover, the non-linearity of the forward model makes the analysis of optimization-based parameter estimation algorithms difficult, while, e.g., sampling-based solutions of the inverse problem may be too time consuming. In [3], the authors considered the inverse problem in the framework of Bayesian particle filtering techniques which avoids the time-consuming forward solutions of the full model over the experimental time window, requiring only local solutions over short time steps. Compared to the particle filtering method with $4800$ particles and $1001$ measurements at $2{\rm Hz}$, the computing times of the proposed dictionary learning method is, given similar hardware, approximately from one half to one third, depending on the hardware. This comparison takes into account the preprocessing time comprising the creation of the dictionary, clustering, extraction of the features and dictionary compression error estimation. We point out that once the preprocessing phase has been completed, solution of the inverse problem is very fast, thus improving the performance by orders of magnitude. While the filtering approach could be made faster, e.g., by resorting to other approximate Bayesian filtering methods, the Ensemble Kalman Fitering (EnKF) or Unscented Kalman Filtering (UKF) in particular being viable candidates, the performance of the dictionary learning technique with precomputed dictionaries remains hard to beat.
@@
The technique proposed in this paper appears useful in the case in which the estimation of parameters must be carried out with limited computational resources, in a relatively short time, or when several estimates must be run in parallel starting with the same dictionary. The obtained estimates appear to be relatively accurate, in fact, the predicted outputs with estimated parameters are indistinguishable from the original data, as Figure 4 suggests. In the numerical experiments with simulated data, the noise was restricted to a coarse truncation error noise. To have a better understanding of the sensitivity, a careful analysis of the measurement noise and inaccuracies should be carried out. A limitation of the proposed algorithm is its reliance on dense sampling of the parameter space which was low-dimensional in the present example. In high-dimensional parameter models, the parameter space sampling requires further thoughts, e.g., by relying multilevel sampling or Monte Carlo sampling with some extra prior information that guides the sampling process.